The municipality of kirkland borrowed 130000 to build a recreation center. The debt principal is to be repaid in eight years, and interest is at 10.5% compounded annually is to be paid annually with end of year payments. To provide for the retirement of the debt the municipal council set up a sinking fund into which equal payments are made at the time of annual interest payments. Interest earned by the fund is 8% compounded annually.



What is annual interest payment on debt.



What is the annual payment into the sinking fund?



What is the total annual cost ( the periodic cost) of the debt?



Compute the book value of the debt after two years.



Compute the interest earned by the fund in year 5.

1
Expert's answer
2022-02-01T15:02:13-0500

1)"P=S(i+\frac{i}{(1+i)^n-1})"

"P=130 000(0.105+\frac{0.105}{(1+0.105)^8-1})=24813.01"


2) the annual payment into the sinking fund is 24813.01


3)"the total annual cost ( the periodic cost) of the debt=24813.01\times8=198504.05"


4)"the book value of the debt after two years=198504.05-24813.01\times2=148 878.04"


5)this is an annuity


"FV=A(\frac{(1+r)^n-1}{r})"


A=24813.01


r=0.08


n=5


"FV=24813.01(\frac{(1+0.08)^5-1}{0.08})=145568"


"145568-24813.01\times5=21502.97"